OpenStudy (jandr):

A captain is twice the age his ship had, when he was as old as his ship is now. Together they are 56 years old. How old is the captain?

OpenStudy (anonymous):

x=captain,y=ship x=2y x-2y=0 x+y=56 slove it and get the age

OpenStudy (jandr):

Re-Read the question x != 2y

OpenStudy (anonymous):

do you have any idea?

OpenStudy (jandr):

re-read the question: together they are 56 years old. 112 + 56 != 56

OpenStudy (anonymous):

let the age of captain be x , and the age of the ship be y , say let some z years before the captain be as old as the ship is now. this means x-z=y; but the captain is old as the age of the ship z years back , ie x=2(y-z); and x+y=56 ,now solve for x =44.8

OpenStudy (jandr):

you got the equasions right, but with them came to the wrong answer. If x=44.8, then y must be 56-44.8=11.2 and z must be x-y=33.6. Z years ago the thip was not built yet.

OpenStudy (jandr):

Apparently this question is too difficult for this audience :-(